OpenStudy (anonymous):

Find the difference of (−4x2 − 7x3 − 2) − (8x3 + 3x − 7x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

(−4x^ − 7x^3 − 2) − (8x^3 + 3x − 7x^2) = (−4x^ − 7x^3 − 2) - 1* (8x^3 + 3x − 7x^2) = (−4x^ − 7x^3 − 2) -1 (8x^3) -1 (3x) -1(− 7x^2) = What comes next, @proverbsm

OpenStudy (anonymous):

just adding this for organization.

OpenStudy (anonymous):

I don't remember the fulll steps, ah.. I know it's using foil right? (first,outside,inside,last) I had this down before , are we now combining like terms? @Directrix

Directrix (directrix):

This is not FOIL. FOIL is a mnemonic some people use for remembering how to multiply a binomial times a binomial. (−4x^2 − 7x^3 − 2) -1 (8x^3) -1 (3x) -1(− 7x^2) = -4x^2 - 7x^3 - 2 - 8x^3 - 3x + 7x^2 = -8x^3 -7x^3 -4x^2 + 7x^2 -3x -2 = ? @proverbsm

OpenStudy (anonymous):

at this point is it combing like terms? @Directrix or do i multiply

Directrix (directrix):

Combing like terms is what you are doing.

OpenStudy (anonymous):

-15x^3 + 3x^2-3x-2

Directrix (directrix):

Yes, that is what I got.

OpenStudy (anonymous):

wow i got that lol

OpenStudy (anonymous):

is that the answer or more to it?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

This is the final answer: -15x^3 + 3x^2-3x-2